Be sure to answer all parts. The reaction of pero
Answer #1

let the rate law be

rate = k[S2082-]^a [I-]^b

now

using exp1 and exp2

rate 2 / rate 1 = ( [I-]2/[I-]1)^b

2.05 x 10-4 / 4.1 x 10-4 = ( 0.017/0.034)^b

b = 1

so

the order with respect to I- is 1

and

simiarly from exp2 and exp3

the order with respect to S2082- is 1

so

the rate law is

rate = k [S2082-] [I-]


now

using exp1

4.1 x 10-4 = k [0.08] [0.034]

k = 0.1507

so

the rate constant is 0.1507 1/Ms